OpenStudy (anonymous):

What is the sum of a 17-term arithmetic sequence where the first term is 6 and the last term is -90?

jimthompson5910 (jim_thompson5910):

use the formula S = n*(a1 + an)/2 where, S = sum n = number of terms a1 = first term an = nth term

OpenStudy (anonymous):

-714 thanks!!! :)

OpenStudy (amistre64):

to get a feel for the formula; consider a more do able set of numbers like: 2+3+4+5+6 , there are 5 numbers entotal if we double this amount and use it in a convenient manner we can work out any arithmetic sum in general; lets reverse the order and add it from underneath 2+3+4+5+6 6+5+4+3+2 ------------ 8+8+8+8+8 notice that each term is equal to the sum of the first and last term: we started with 5 numbers: 5(8) = 5(2+6) = 40 but we doubled the value to begin with so we want to halve it to correct it back 40/2 = 20 in general, we can add the first and last terms; multiply it by how many terms there are, and divide the results in half

jimthompson5910 (jim_thompson5910):

-714 is correct

OpenStudy (anonymous):

Thank you guys! :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (amistre64):

good luck